log(3+x) - log(x-3)=2log2
Answer by stanbon(75887)

log(3+x) - log(x-3)=2log2
-------
log[(3+x)/(x-3)] = log(4)
-------
(3+x)/(x-3) = 4
------
3+x = 4x - 12
-----
3x = 15
x = 5
----------------
Cheers,
Stan H.
-------------